# Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions

### v3

@@ 3.5. Efficiency Study (For RQ3)
Figure 5. Generalization across Datasets, LLAMA-DES-H. The horizontal axis represents the training set, and the vertical axis represents the test set.
Figure 6. Performance of LLAMA-DES2 (GPT) versus the number of sampled triplets.
@@ 4.1. Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. And (Li et al., 2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning. More recently, KG-LLAMA (Yao et al., 2024) is the first to adopt a decoder-only LLM architecture for KGC by formulating the task as question answering through an instruction-tuning-based prompt, followed by parameter-efficient fine-tuning on LLaMA. Notably, KG-LLAMA leverages only LLaMA’s own capabilities without incorporating additional hybrid approaches, which aligns with the core principle of our proposed framework.
@@ 4.2. Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by (Saier et al., 2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. (Ramrakhiyani et al., 2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, (Jin et al., 2024) used probing to measure the abilities of different LLM layers to learn various levels of concepts. In addition, (Liu et al., 2024) investigated whether a universal truthfulness hyperplane exists within LLMs by employing probing methods.

### v4

@@ 3.6. Additional Analysis
Figure 8. Generalization across Datasets, LLAMA-MLP-DES (NG). The horizontal axis represents the training set, and the vertical axis represents the test set.
@@ 4.1. Knowledge Graph Completion
Knowledge graphs are inherently incomplete, necessitating effective approaches for knowledge graph completion (Hogan et al., 2021). Prevailing KGC methodologies generally fall into two categories: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural patterns of knowledge graphs. Representative methods such as TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), ConvE (Dettmers et al., 2018) and RotatE (Sun et al., 2019) primarily use scoring functions defined in a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. Representative methods include KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, reformulating KGC as a sequence classification problem solvable via encoder-only language models. Meanwhile, KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) leverage encoder-decoder language models to perform KGC in a generative manner. Recently, decoder-only large language models have shown great promise in KGC tasks. KoPA (Zhang et al., 2024) integrates structural embeddings with LLMs to jointly leverage the structural information of KGs and their powerful contextual modeling capabilities. More recently, KG-LLAMA (Yao et al., 2025) frames KGC as an instruction-based question-answering task, applying parameter-efficient fine-tuning techniques to LLMs. Unlike hybrid approaches that combine LLMs with external structural embeddings, KG-LLAMA solely relies on the intrinsic capabilities of LLaMA, aligning closely with the core motivation of our proposed framework.
@@ 4.2. Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual knowledge. As demonstrated by (Saier et al., 2023), LLMs serve as implicit knowledge bases, encoding extensive world knowledge within their parameters. This intrinsic capability positions LLMs as promising tools for knowledge-intensive tasks such as knowledge graph completion, by effectively leveraging their embedded factual representations. To examine the extent and quality of factual knowledge encoded in LLMs, recent works have employed probing methods. For instance, (Ramrakhiyani et al., 2023) conducted evaluations across multiple LLMs to assess their knowledge of world geography using probing. Similarly, (Jin et al., 2024) analyzed the conceptual abstraction capabilities of different model layers, revealing hierarchical encoding patterns within the intermediate representations. Furthermore, (Liu et al., 2024) investigated the potential existence of a universal truthfulness hyperplane in LLMs via probing. Notably, recent probing analyses (Skean et al., 2025, [n. d.]) have indicated that intermediate layers often yield more informative representations for downstream tasks than the final layers. These findings align closely with our approach of leveraging intermediate representations from LLMs to achieve effective and efficient KGC.